# Universality Diagram of Phase Transitions in Long-range Statistical Systems

### v1

@@ VII Conclusion
We integrate recent large-scale numerical simulations, rigorous mathematical results, and known field-theoretic limits (SR, GF, and CG) into a single, cohesive picture across these models. A central conjecture of our work is the division of the nonclassical LR regime (LR-LD) into two distinct phases, separated by the boundary $\sigma=1$. Based on combined evidence and insights from LR-SRW, we propose that in the regime $\sigma\leq 1$ (LR-LD-B), the anomalous dimension retains its LR-Gaussian prediction $\eta=2-\sigma$, even as the thermal exponent $y_{t}$ deviates. In this regime, the hyperballistic behavior of LR-SRW implies strong nonlocality in the field-theoretical description of the model. On this basis, we conjecture that, for $\sigma<1$, the critical behaviors in the LR-percolation and LR-O($n$) models do not exhibit conformal invariance and cannot be described by conformal field theory. In contrast, for the $1<\sigma\leq 2$ regime (LR-LD-A), we posit that all critical exponents, including $\eta$, deviate from their mean-field values.
@@
Our perspective of long-range criticality in the $(\sigma,d)$ plane of these foundational models shall provide a conceptual framework for guiding future analytical and numerical investigations. We emphasize that, while various aspects of our picture are consistent with numerical simulations, mathematical results, and are based on physically insightful arguments, the universality diagrams in Figs. 13 are conjectured and remain to be confirmed or falsified in the future. In particular, the boundaries, $\sigma=d$ for $d<2$ and $\sigma=1$ dividing the LR-LD regime, are proposed for the first time. In addition, the boundary $\sigma=2$ contradicts Sak’s criterion, but is supported by a series of numerical simulations [picco2012, xiao_two-dimensional_2024, yao2025nonclassicalregimetwodimensionallongrange, liu_two-dimensional_2025]. For the FK-Ising model in the mean-field regimes, including SR-HD-A, SR-HD-B, LD-HD-A, and LD-HD-B, a set of critical exponents has been conjectured, and the nature of logarithmic corrections at marginal boundaries remains to be investigated.

### v2

@@ VII Conclusion
We integrate recent large-scale numerical simulations, rigorous mathematical results, and known field-theoretic limits (SR, GF, and CG) into a single, cohesive picture across these models. A central conjecture of our work is the division of the nonclassical LR regime (LR-LD) into two distinct phases, separated by the boundary $\sigma=1$. Based on combined evidence and insights from LR-SRW, we propose that in the regime $0<\sigma\leq 1$ (LR-LD-B), the anomalous dimension retains its LR-Gaussian prediction $\eta=2-\sigma$, even as the thermal exponent $y_{t}$ deviates. In this regime, the hyperballistic behavior of LR-SRW implies strong nonlocality in the field-theoretical description of the model. On this basis, we conjecture that, for $0<\sigma<1$, the critical behaviors in the LR-percolation and LR-O($n$) models do not exhibit conformal invariance and cannot be described by conformal field theory. In contrast, for the $1<\sigma\leq 2$ regime (LR-LD-A), we posit that all critical exponents, including $\eta$, deviate from their mean-field values.
@@
Our perspective of long-range criticality in the $(\sigma,d)$ plane of these foundational models shall provide a conceptual framework for guiding future analytical and numerical investigations. We emphasize that, while various aspects of our picture are consistent with numerical simulations, mathematical results, and are based on physically insightful arguments, the universality diagrams in Figs. 13 are conjectured and remain to be confirmed or falsified in the future. In particular, the boundaries, $\sigma=d$ for $d<2$ and $\sigma=1$ dividing the LR-LD regime, are proposed for the first time. In addition, the boundary $\sigma=2$ contradicts Sak’s criterion, but is supported by a series of numerical simulations [44, 56, 58, 38]. For the FK-Ising model in the mean-field regimes, including SR-HD-A, SR-HD-B, LD-HD-A, and LD-HD-B, a set of critical exponents has been conjectured, and the nature of logarithmic corrections at marginal boundaries remains to be investigated.